OpenStudy (anonymous):

I WILL FAN AND MEDAL!!! HELP? The perimeter of a rectangle is four more than the area of the rectangle. Find the area of the rectangle, if its width is 1 cm.

OpenStudy (welshfella):

let the length be x cm then the area = x*1 = x cm^2 the perimeter = 2 + 2x and the perimeter is 4 more than the area so 2x + 2 = x + 4 solve for x to find the area

OpenStudy (anonymous):

give me a sec. to figure this one mate

OpenStudy (anonymous):

2?